OpenStudy (anonymous):

Use the function: f(x) = x3 + 2 Which statement is true? Select one: a. The equation has an inverse as it has an answer and the Horizontal Line Test proves it has an inverse. b. The equation has an answer, but the Horizontal Line Test proves it does not have an answer. c. The equation does not have an answer and has no inverse. d. The equation passes the Horizontal Line Test but does not have an answer.

OpenStudy (anonymous):

@UnkleRhaukus @iGreen

OpenStudy (anonymous):

@texaschic101 @undeadknight26 @jhonyy9

OpenStudy (anonymous):

you are not getting a response because the question makes no sense

OpenStudy (anonymous):

\[ f(x) = x^3 + 2\] what does "The equation has an answer" mean ?

OpenStudy (anonymous):

if you solve it you will come up with an answer?

OpenStudy (anonymous):

no it means nothing there is no such thing as "solving" \(f(x)=x^3+2\) can you post the exact question? maybe a screenshot?

OpenStudy (anonymous):

that is the exact question

OpenStudy (anonymous):

really? where does it come from ? who wrote it?

OpenStudy (anonymous):

I take online classes

OpenStudy (anonymous):

\[f(x)=x^3+2\] the FUNCTION does have an inverse, it is \[f^{-1}(x)=\sqrt[3]{x-2}\] the "horizontal line test" will prove that it has an inverse so maybe A is the correct answer the rest is gibberish would still like to know where the question came from

OpenStudy (anonymous):

so I should go with A